# F-LMM: Grounding Frozen Large Multimodal Models

### v1

@@ 4 Experiments
Grounding Frozen General-Purpose LMMs by F-LMM (Ours)
DeepseekVL-1.3B [42] 1307/225 64.6 34.8 51.1 75.0 62.8 68.2 64.9 63.4 68.3
MGM-2B [32] 1341/312 59.8 31.1 65.9 75.0 63.7 67.3 65.6 64.4 68.4
LLaVA-1.5-7B [37] 1511/348 64.3 30.5 69.0 75.2 63.7 67.1 64.8 63.4 68.2
HPT-Air-6B [62] 1010/ 258 69.8 31.3 59.2 74.3 64.0 67.5 65.5 64.0 68.8
HPT-Air-1.5-8B [62] 1476/308 75.2 36.3 62.1 76.3 64.5 68.5 65.4 64.1 68.5
MGM-7B [32] 1523/316 69.3 40.8 75.8 75.7 64.8 68.3 66.3 65.3 68.6
DeepseekVL-7B [42] 1468/298 73.2 41.5 77.8 76.1 66.4 70.1 65.7 64.5 68.5
LLaVA-1.6-7B [38] 1519/322 68.1 44.1 72.3 75.8 65.8 70.1 66.3 65.1 69.0
LLaVA-1.6-M-7B [38] 1501/324 69.5 47.8 71.7 75.7 66.5 70.1 66.5 65.4 69.1
MGM-HD-7B [32] 1546/319 65.8 41.3 74.0 76.1 65.2 68.5 66.7 65.6 69.1
4.1 Implementation Details
Model Architectures. We build F-LMM on several open-sourced LMMs, including LLaVA-1.5 [37], LLaVA-Next [38], MiniGemini [32], DeepseekVL [42] and HPT-Air [62]. The main experiment covers 10 LMMs with model sizes ranging from 1.3B to 8B. We employ a lightweight 3-stage U-Net [59] as the CNN-based mask decoder to transform segmentation priors from frozen LLMs. The U-Net architecture features an encoder-decoder structure with skip connections, wherein feature maps are downsampled in the encoder and upsampled in the decoder. Please check Sec A.2 of the appendix for more details on the mask decoder. As for the SAM-based mask refiner, we choose SAM ViT-L [26] that balances cost and performance well.
Model Training. F-LMM is implemented on XTuner [13]. We train F-LMM on RefCOCO(+/g) [23, 44] and PNG [17] datasets with about 190k data samples on a single machine with 8 NVIDIA A800-40G GPUs, which costs about 20 hours for each round of model training. We set the batch size to 8 and train models for 8 epochs, with gradient clipping at a max norm of 1.0. The AdamW [41] optimiser is used with a learning rate of 1e-4, a weight decay of 0.01, and betas as (0.9, 0.999). We choose a warm-up ratio of 0.03 at the beginning of training to stabilise model optimisation.
4.2 Main Evaluation
Our main evaluations cover both the conversational and grounding ability of LMMs. We summarise the evaluation results of grounding LMMs in Table 1. Please refer to Sec A.1 in the Appendix for more detailed results.
Comparisons with Existing Methods. We compare F-LMM with existing grounding LMMs. As shown in Table 1, our F-LMM provides the best balance with conversational and grounding abilities among compared methods. On the question-answering benchmarks, existing grounding LMMs obtain zero or near-zero scores on MMBench and MME while lagging significantly behind general-purpose LMMs on MMVet and LLaVA-In-the-Wild benchmarks, indicating compromised instruction-following ability and weakened general knowledge comprehension. On the RES and PNG benchmarks, our F-LMM achieves comparable results despite not having the parameters of LMMs fine-tuned for grounding purposes. Compared with standard segmentation models, F-LMM outperforms all these specially designed models on both RES and PNG benchmarks.
@@ 4.2 Main Evaluation
F-LMM (Ours)
DeepseekVL-1.3B [42] 30.4 58.2 69.7 23.9 20.0 31.0 87.4 86.6
DeepseekVL-1.3B [42] 38.6 62.2 75.0 31.8 31.6 34.4 88.3 88.1
DeepseekVL-7B [42] 43.2 63.5 74.5 32.0 28.4 43.2 87.0 86.0
DeepseekVL-7B [42] 53.8 67.9 78.4 42.3 44.1 49.1 88.0 87.7
4.3 Unleashing Visual Chain-of-Thought Reasoning
Considering our F-LMM can ground objects without losing any instruction-following ability, we further study whether visual chain-of-thought (CoT) reasoning can be elicited in our model. In human-AI conversations that involve Visual CoT, an LMM first localises the region/object relevant to the human’s question and then generates the final answer by zooming in on the question-related region. Here, we use DeepseekVL-1.3B and DeepseekVL-7B models that support multi-visual inputs and evaluate on the VisCoT benchmark [60]. As shown in Table 2, our models achieve remarkable performance gains when prompted in a visual CoT manner. It is noticeable that our F-LMM even outperforms VisCoT-7B [60] that has been well-tuned on the training set of VisCoT data [60]. Furthermore, we perform visual CoT reasoning on the object hallucination benchmark POPE [34] and observe significant performance gain in resisting object hallucinations. Thanks to the combination of excellent grounding and instruction-following abilities, our F-LMMs have the potential to perform complex visual perception and reasoning. For more details of visual CoT reasoning on the VisCoT benchmark [60], please refer to Figure 6.
Table 3: Ablation study of F-LMM (based on DeepseekVL-1.3B) on the PNG benchmark. The blue colour indicates our default design choice.
((a)) Plain CNN v.s. U-Net
# Mask PNG
Decoder All Thing Stuff
1 Plain CNN 64.5 63.1 67.8
2 U-Net 64.9 63.4 68.3
((b)) Merge Type
# Merge PNG
Type All Thing Stuff
1 Max 64.6 63.1 68.0
2 Average 64.9 63.4 68.3
((c)) Input Normalisation
# Normalize PNG
Input All Thing Stuff
1 64.5 63.0 68.0
2 64.9 63.4 68.3
((d)) Input Size
# Input PNG
Size All Thing Stuff
1 32 64.2 62.8 67.6
2 64 64.9 63.4 68.3
3 128 65.0 63.5 68.4
((e)) SAM Variants
# SAM PNG
Variant All Thing Stuff
1 ViT-B 63.0 61.4 66.8
2 ViT-L 64.9 63.4 68.3
3 ViT-H 65.0 63.5 68.3
((f)) Prompts for Mask Refiner
# Prompts PNG
@@ 4.3 Unleashing Visual Chain-of-Thought Reasoning
1 63.4 62.0 69.8
2 63.7 62.2 67.1
3 64.9 63.4 68.3
Table 4: Analysis of original visual grounding ability in LMMs by discarding object-centred data samples during training.
# Model VisualGenome RES PNG
data RefCOCO RefCOCO+ RefCOCOg All Thing Stuff
1 LLaVA-1.5-7B [37] 75.2 63.7 67.1 64.8 63.4 68.2
2 LLaVA-1.5-7B [37] 73.2 60.8 65.1 64.5 63.0 67.9
4.4 Ablation Study
We investigate the effects of design choices of F-LMM. All the ablation studies are conducted on the PNG benchmark (‘All’) using the smallest LMM, DeepseekVL-1.3B [42].
Mask Decoder. We consider two architecture variants of the mask decoder: a U-Net [59] involves several downsampling and upsampling operations and a plain CNN without variations in the resolution of feature maps. To ensure a fair comparison, we keep both two architectures having approximately the same number of parameters (i.e., 8M). As shown in Table 3, the U-Net outperforms the plain CNN by a 0.4 margin, substantiating the importance of multi-scale structure in building segmentation models. Then, we study the effectiveness of normalizing word-image attention maps, which ensures all attention scores sum up to 1.0. By eliminating the influence of varying sequence length on the magnitude of attention scores, the normalisation of the attention map can provide a performance increase of 0.4, as shown in Table 3. In addition, we also find that averaging attention maps of a multiple-word object achieves better performance than max operation by a margin of 0.3. Finally, we study the input size of the mask decoder in Table 3, including $32\times 32$, $64\times 64$, and $128\times 128$, observing the performance increases with the resolution of inputs. By default, we choose $64\times 64$ to balance cost and performance.
Mask Refiner. We consider three SAM [26] mode variants, i.e., ViT-B(ase), ViT-L(arge) and ViT-H(uge) as mask refiners. As shown in Table 3, the performance grows with model sizes. For a good trade-off between cost and performance, we select ViT-L as the default mask refiner. To analyze the effect of different prompts on the mask refiner, we start by only using coarse mask logits from the CNN-based mask decoder. Then, we gradually add the bounding box of the coarse masks and text embeddings as prompts. As shown in Table 3, the performance increases when we stack up more prompts for mask refinement. It is evident that utilizing the coarse mask logits only can already yield considerable performances, verifying that the geometric and spatial cues from the transformers’ attention mechanism are effective for visual grounding.
4.5 Analysis and Visualisation
Model Chat Ground
@@ 4.5 Analysis and Visualisation
Figure 5: The left table shows the average ranks of each LMM on question-answering (‘Chat’) and grounding benchmarks (‘Ground’). The dashed line in the right figure is the linear fit of the rank data points, indicating a positive correlation between abilities to chat and ground.
Figure 6: Examples of visual chain-of-thought reasoning. The model used is DeepseekVL-1.3B and the samples are taken from the test set of VisCoT data [60]. The LMM is first prompted to think about the question-related object, which is grounded by the mask head of F-LMM. The region of the question-related object is cropped and fed to the LMM to help answer the question.
The Origin of Visual Grounding Ability. One might question whether the visual grounding ability in LMMs originates from object-centred training samples. For example, a considerable part of the LLaVA-Instruct data [37] is from VisualGenome [27] dataset in which question-answer pairs are based on specific regions or objects. Regarding this concern, we discard the data samples that are taken from the VisualGenome and train an LLaVA-1.5-7B model on the rest of the LLaVA-Instruct data. As shown in Table 4(#2), we can still obtain competitive performance on the grounding benchmarks. Therefore, we believe the built-in grounding ability in LLMs’ attention mechanism can be learned even if the LMMs are trained only on image-level data samples.
@@
Better Chatting Means Better Grounding? We study the correlation between performance on the question-answering benchmarks and the grounding benchmarks. For the 10 models reported in Table 1, we calculate their average ranks in each benchmark category (the lower the rank, the better the performance). In Figure 5, we visualise the two types of ranks as 2D coordinates, i.e., (Chat Rank, Ground Rank), and linearly fit these rank data points. As indicated by the blue dashed line, frozen LMMs with better conversational ability would serve as better backbones for grounding. We also observe that larger LMMs tend to be better at both conversation and grounding, and LMMs with larger input resolution (e.g., LLaVA-1.6 and MGM-HD) can handle both tasks better.
Visualisation of Visual CoT. In Figure 6, we show examples of visual chain-of-thought reasoning (Visual CoT), in which the LMM is first prompted to answer "What object is the most relevant to the question?". Then, the mask head of F-LMM grounds the LMM’s answer about the relevant object by generating a segmentation mask, the bounding box of which is used to crop the object region from the original image. Finally, the cropped image region is fed into the LMM to obtain the final answer. As shown in Figure 6, the Visual CoT powered by the LMM’s grounding ability is helpful when the LMM needs to focus on the question-related regions for better visual perception and reasoning.

### v2

@@ 3.1 Segmentation Priors from Frozen LMM
F-LMM (DeepSeekVL-1.3B [44]) 1307/225 64.6 34.8 51.1 75.0 62.8 68.2 64.9 63.4 68.3
F-LMM (MGM-2B [34]) 1341/312 59.8 31.1 65.9 75.0 63.7 67.3 65.6 64.4 68.4
F-LMM (LLaVA-1.5-7B [38]) 1511/348 64.3 30.5 69.0 75.2 63.7 67.1 64.8 63.4 68.2
F-LMM (HPT-Air-6B [65]) 1010/ 258 69.8 31.3 59.2 74.3 64.0 67.5 65.5 64.0 68.8
F-LMM (HPT-Air-1.5-8B [65]) 1476/308 75.2 36.3 62.1 76.3 64.5 68.5 65.4 64.1 68.5
F-LMM (MGM-7B [34]) 1523/316 69.3 40.8 75.8 75.7 64.8 68.3 66.3 65.3 68.6
F-LMM (DeepSeekVL-7B [44]) 1468/298 73.2 41.5 77.8 76.1 66.4 70.1 65.7 64.5 68.5
F-LMM (LLaVA-1.6-7B [40]) 1519/322 68.1 44.1 72.3 75.8 65.8 70.1 66.3 65.1 69.0
F-LMM (LLaVA-1.6-M-7B [40]) 1501/324 69.5 47.8 71.7 75.7 66.5 70.1 66.5 65.4 69.1
F-LMM (MGM-HD-7B [34]) 1546/319 65.8 41.3 74.0 76.1 65.2 68.5 66.7 65.6 69.1
Segmentation Priors in Self-Attention. The vision-language sequence is mainly processed by causal self-attentions [68, 56] in the LLM, including inner product and weighted-sum operations. Specifically, for a word token with position index $i$ in the vision-language sequence ${\mathbf{Z}}$ , its embedding ${\bm{z}}^{i}$ is updated by the weighted sum of the first $i$ embeddings: $\hat{{\bm{z}}}^{i}=\texttt{SoftMax}(\frac{{\bm{z}}^{i}\cdot{\mathbf{Z}}[:i]}{d})\cdot{\mathbf{Z}}[:i]$, where $\texttt{SoftMax}(\frac{{\bm{z}}^{i}\cdot{\mathbf{Z}}[:i]}{d})$ is the attention weights. Here, we omit the residual layers and feedforward layers for brevity.
$\bm{a}^{i}=\texttt{Unflatten}(\texttt{SoftMax}(\frac{{\bm{z}}^{i}\cdot{\mathbf{Z}}[:i]}{d})[:hw])\in{\mathbb{R}}^{h\times w},$
where Unflatten restores the 2-D spatial structure from the 1-D sequence to form an attention map. In Figure 4, we visualise such word-image attention maps from various transformers layers and attention heads in an LMM (i.e., DeepSeekVL-7B [44]). The objects’ shape and location can be observed in word-image attention maps of certain layers or heads. The visibility is further enhanced when we stack the attention maps from all layers and heads and perform K-Means clustering. It can be observed that the attention maps offer meaningful segmentation priors with spatial and geometric cues for grounding objects visually.
3.2 Visual Grounding with Mask Head
Mask Decoder. The mask decoder $f_{d}$ is a 2-D CNN model that transforms the word-image attention maps of grounded objects into mask logits, which is instantiated by a 3-stage U-Net [62]. Please refer to the supplemental material for details on the mask decoder. The extraction of word-image attention map $\bm{a}^{i}$ for a word token with position index $i$ is illustrated in Eq. 3.1 and Figure 3. For an object described by multiple words, we merge their corresponding word-image attention maps to a single attention map $\bm{a}$ via element-wise average or max operation. The attention map $\bm{a}$ is further normalised as $\bm{a}/\mathrm{sum}(\bm{a})$ so that all elements sum to 1. Considering $M$ layers and $N$ attention heads, we stack the $MN$ attention maps as $\bm{A}\in{\mathbb{R}}^{MN\times h\times w}$, which forms the input to a mask decoder. Given the importance of high input resolution for segmentation models, we upsample the stacked attention maps $\bm{A}$ to $h^{\prime}\times w^{\prime}$ by bilinear interpolation before feeding it to a mask decoder, where $h^{\prime}>h$ and $w^{\prime}>w$. In practice, we set $h^{\prime}=w^{\prime}=64$. Then, the mask decoder maps $\bm{A}$ into mask logits: ${\mathbf{M}}_{\mathrm{logits}}=f_{d}(\bm{A})$. We derive the corresponding binary mask via ${\mathbf{M}}_{\mathrm{pred}}={\mathbf{M}}_{\mathrm{logits}}>0$. During training, the mask decoder is optimised with BCE and DICE losses [64].
Table 2: Reasoning Segmentation.
Model Val Test
- Short Long All
X-Decoder [88] 22.6 20.4 22.2 21.7
SEEM [89] 25.5 20.1 25.6 24.3
GroundingSAM [41] 26.0 17.8 22.4 21.3
OVSeg [35] 28.5 18.0 28.7 26.1
LISA [28] 44.4 37.6 36.6 36.8
F-LMM 46.7 36.9 49.1 46.2
Table 3: Grounded Conversation Generation (GCG). M. stands for METEOR.
Model GCG Val Test
Training M. mIoU Recall M. mIoU Recall
LISA [28] 13.0 62.0 36.3 12.9 61.7 35.5
OMG-LLaVA [82] 14.9 65.5 - 14.5 64.7 -
GLaMM [60] 16.2 66.3 41.8 15.8 65.6 40.8
BuboGPT [86] 17.2 54.0 29.4 17.1 54.1 27.0
KOSMOS-2 [52] 16.1 55.6 28.3 15.8 56.8 29.0
F-LMM 17.6 63.5 42.0 17.4 63.6 38.6
@@ 3.2 Visual Grounding with Mask Head
F-LMM 43.2 63.5 74.5 32.0 28.4 43.2 87.0 86.0
F-LMM 53.8 67.9 78.4 42.3 44.1 49.1 88.0 87.7
3.3 Keyword Selector for Grounded Conversation
In grounded conversation [60] with interleaved segmentation masks and words, existing grounding LMMs [60, 61] expand LLMs’ vocabularies with special tokens that indicate the start and end of grounding targets, which is infeasible in F-LMM given its ‘frozen‘ nature. A common practice to discover visual objects in text sequences offline is using external tools such as SpaCy [20], which parse all nouns from a sentence including unwanted non-object words. Instead of adopting such offline tools that produce noisy results, we automate generating interleaved words and masks by training a linear layer to directly predict whether a word is to be grounded or not.
Specifically, the linear layer (keyword selector) is placed on top of the LLM’s transformer layers, projecting $d$-dimension hidden state vectors into one-dimension scores, followed by a sigmoid function that normalizes the scores to $[0,1]$. During training, the score prediction is supervised by a BCE loss. During inference, word tokens with scores exceeding a threshold $\lambda$ are regarded as positive for visual grounding. In practice, we set $\lambda=0.3$. Adjacent positive word tokens are grouped to indicate a single visual object. After the word tokens of visual objects are selected, the corresponding attention maps and text embeddings are fed to the mask head for visual grounding.
4 Experiments
4.1 Implementation Details
Model Architectures. We build F-LMM on several open-sourced LMMs, including LLaVA-1.5 [38], LLaVA-Next [40], MiniGemini [34], DeepSeekVL [44] and HPT-Air [65]. The main experiment covers 10 LMMs with model sizes ranging from 1.3B to 8B. We employ a lightweight 3-stage U-Net [62] as the mask decoder to transform segmentation priors from frozen LLMs. The U-Net architecture features an encoder-decoder structure with skip connections, wherein feature maps are downsampled in the encoder and upsampled in the decoder. Please check the supplementary material for more details on the mask decoder. As for the SAM-based mask refiner, we choose SAM ViT-L [26] that balances cost and performance well.
Model Training. We train F-LMM on RefCOCO(+/g) [24, 46] and PNG [16] datasets with about 190k data samples on a single machine with 8 NVIDIA A800-40G GPUs, which costs about 20 hours for each round of model training. We set the batch size to 8 and train models for 8 epochs, with gradient clipping at a max norm of 1.0. The AdamW [43] optimiser is used with a learning rate of 1e-4, a weight decay of 0.01, and betas as (0.9, 0.999). We choose a warm-up ratio of 0.03 to stabilise model optimisation.
4.2 Standard QA and Grounding Tasks
For a comprehensive study of LMMs’ conversational and grounding capabilities, we first evaluate models under standard question-answering and grounding benchmarks separately. We summarise the evaluation results of grounding LMMs in Table 1. Please refer to the supplementary material for more detailed results.
Comparisons with Existing Methods. We compare F-LMM with existing grounding LMMs. As shown in Table 1, our F-LMM provides the best balance with conversational and grounding abilities among compared methods. On the question-answering benchmarks, existing grounding LMMs obtain zero or near-zero scores on MMBench and MME while lagging significantly behind general-purpose LMMs on MMVet and LLaVA-In-the-Wild benchmarks, indicating compromised instruction-following ability and weakened general knowledge comprehension. On the RES and PNG benchmarks, our F-LMM achieves comparable results despite not having the parameters of LMMs fine-tuned for grounding purposes.
Figure 5: Ablation study of the mask decoder.
4.3 Complex Scenarios
In this section, we evaluate grounding LMMs under more complex scenarios that typically require the LMMs to perform both reasoning and segmentation. The base LMM we use is DeepSeekVL-7B [44], considering its flexibility in supporting both single and multiple image inputs. And the size of models compared in this section is also 7B.
Reasoning Segmentation is proposed in LISA [28] that requires a model to infer what object to segment from common-sense knowledge or via logical reasoning. The evaluation results are provided in Table 3 and the metric is gIoU. F-LMM can effectively perform reasoning segmentation even though it is not trained on such type of data. It is remarkable that F-LMM significantly outperforms existing models on the subset of long sentences, reflecting the advantage of F-LMM in handling complex contexts.
Grounded Conversation Generation (GCG) is proposed in GLaMM [60], which requires a model to generate interleaved segmentation masks and texts. For performance evaluation, METEOR (M.) and mIoU are used to measure the quality of generated texts and masks, respectively. In addition, we report the recall of object masks. As shown in Table 3, F-LMM exhibits the best zero-shot performance while being comparable with models fine-tuned on GCG training dataset.
Table 5: Ablation study of the mask refiner on PNG benchmark.
(a) Effects of Mask Refiner
Mask Refiner PNG
@@ 4.3 Complex Scenarios
50.8 48.6 55.9
63.4 62.0 66.8
63.7 62.2 67.1
64.9 63.4 68.3
(b) SAM Variants
SAM PNG
All Thing Stuff
ViT-B 63.0 61.4 66.8
ViT-L 64.9 63.4 68.3
ViT-H 65.0 63.5 68.3
Table 6: Ablation study of keyword selection.
Method F1 Recall Precision
SpaCy Parser [20] 57.8 97.3 41.1
Linear Keyword Selector 82.8 96.6 72.5
Model Chat $\downarrow$ Ground $\downarrow$
@@
Figure 6: The relevance between an LMM’s chat ability and the grounding ability of the F-LMM built on it. The left table shows the average ranks of each LMM on question-answering (‘Chat’) and grounding benchmarks (‘Ground’). $\downarrow$ means the lower the better. The dashed line in the right figure is the linear fit of the rank data points, indicating a positive correlation between abilities to chat and ground.
Visual Chain-of-Thought Reasoning. In human-AI conversations that involve Visual Chain-of-Thought Reasoning (VisCoT) [63], an LMM first localises the region/object relevant to the human’s question and then generates the final answer by zooming in on the question-related region. Here we evaluate F-LMM and VisCoT-7B on the VisCoT benchmark [63]. As shown in Table 4, F-LMM achieves remarkable performance gains when prompted in the VisCoT manner. It is noticeable that F-LMM even outperforms VisCoT-7B [63] that has been well-tuned on the VisCoT training data [63]. Furthermore, we perform VisCoT on the object hallucination benchmark POPE [33] and observe significant performance gain in resisting object hallucinations.
4.4 Ablation Study
We investigate the effects of design choices of F-LMM. All the ablation studies are conducted on the PNG dataset and we use DeepSeekVL-1.3B [44] unless otherwise stated.
Mask Decoder. We summarise our analyses of the mask decoder in Figure 5. Note that the mask refiner is not involved in this part. We first consider attention maps from different transformer layers, i.e., early ($6^{\text{th}}$), mid ($12^{\text{th}}$) and late ($24^{\text{th}}$) layers. As shown in Figure 5(a), attention maps from late layers perform the worst, conforming to prior studies [21] indicating that deeper transformer layers tend to focus on abstract concepts instead of visual details. And using attention maps from all layers achieves the best performance. Next, we study how to merge the word-image attention maps of multi-word objects, as shown in Figure 5(b). We find that the average operation outperforms the max operation by a margin of 0.7. In Figure 5(c), we show that normalizing the inputs to the mask decoder provides a 0.5 performance gain. Finally, we experiment with different input sizes for the mask decoder. As shown in Figure 5(d), using $64\times 64$ yields the best performance-cost trade-off.
Mask Refiner. We study the effects of the mask refiner on segmentation in Table 5. The performance of using only the mask decoder is shown in the first row of Table 5. With the masks fed to the mask refiner, we observe a significant 12.6 performance gain on the PNG benchmark. Adding box and text prompts to the mask refiner further improves the performance by 0.3 and 1.2. Then we experiment with different SAM [26] model variants, i.e., ViT-B(ase), ViT-L(arge) and ViT-H(uge) as mask refiners. As shown in Table 5, the performance grows with model sizes. For a good trade-off between cost and performance, we select the ViT-L variant of SAM as the default mask refiner.
Keyword Selector. We analyse the keyword selector, implemented as a linear layer, on the PNG dataset, using the F1 score as the main metric. As shown in Table 6, our keyword selector achieves significantly higher F1 scores than the external SpaCy tool [20]. We also report recall and precision scores. Our keyword selector achieves comparable recall while being much more precise compared to SpaCy, which enumerates all nouns in a sentence.
4.5 Analysis & Visualisation
@@
Scalability: Does Better Chatting Lead to Better Grounding? We study the relevance between an LMM’s chat ability and the grounding ability of the F-LMM built on it. Specifically, we examine the correlation between performance on the question-answering and grounding benchmarks. For the ten models reported in Table 1, we calculate their average ranks in each benchmark category. In Figure 6, we plot these ranks as 2D coordinates, i.e., (Chat Rank, Ground Rank), and apply a linear fit to the data points. As indicated by the blue dashed line, frozen LMMs with stronger conversational ability can serve as better backbones for grounding. We also observe that larger LMMs generally excel in both conversation and grounding tasks, and LMMs with larger input resolutions (e.g., LLaVA-1.6 and MGM-HD) can handle both tasks better.
Figure 7: Visualisations of KMeans clustering on attention maps and segmentation results from the CNN-based mask decoder and the SAM-based mask refiner.
From Attention Maps to Segmentation Masks. We visualise the word-image attention maps by applying KMeans clustering to the stacked attention maps that are collected from all transformer layers and attention heads. The attention maps of multiple-word objects are merged by element-wise average. As shown in Figure 7, we observe that the pixels of objects are roughly clustered together (top-left). With the CNN-Based mask decoder, the attention weights are mapped to 2D binary masks (top-right), which are then further optimised by the SAM-based mask refiner (bottom-left). The model used is DeepSeekVL-1.3B [44].